Write wavefunction for hydrogen vatom for ground state.
1
Expert's answer
2020-11-02T10:53:05-0500

Solution

Wave function for hydrogen atom can be written as

n=1 ,l=0, m=0


"\psi_{100}=\sqrt\frac{1}{\pi a_0^3} e^{-\frac{r}{2a_0}}"